OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the equation 15x - 5y = 30. x-intercept = (-2, 0); y-intercept = (0, 6) x-intercept = (2, 0); y-intercept = (0, -6) x-intercept = (0, -6); y-intercept = (2, 0) x-intercept = (0, 6); y-intercept = (-2, 0) @eSpeX ?

OpenStudy (anonymous):

x intercept is when y=0 y intercept is when x=0 so to find x intecept, substitute y=0 then solve for x same thing for y intercept

OpenStudy (anonymous):

Can you show me? That part is always confusing.

OpenStudy (anonymous):

15x - 5y = 30 to find y intecept x=0 15(0) - 5y = 30 solve for y then the y intecept will be the point (0, yo) where yo is the value you solved for

jimthompson5910 (jim_thompson5910):

Finding the y-intercept 15x - 5y = 30 15(0) - 5y = 30 ... plug in x = 0 0 - 5y = 30 -5y = 30 y = 30/(-5) y = -6 So the y-intercept is the point (0, -6)

jimthompson5910 (jim_thompson5910):

Finding the x-intercept 15x - 5y = 30 15x - 5(0) = 30 ... plug in y = 0 15x - 0 = 30 15x = 30 x = 30/15 x = 2 So the x-intercept is (2, 0)

OpenStudy (anonymous):

Thanks @jim_thompson5910 that makes so much more sense :)

jimthompson5910 (jim_thompson5910):

you're welcome